Existence and approximation of radial solutions of semilinear elliptic equations in an annulus

B. Vrdoljak, Split, Croatia

Keywords: Semilinear elliptic equations, radial solutions, existence, approximation AMS Subject classification: 35B30, 35B40, 35J60

Abstract: This paper deals with the existence, behavior and approximation of positive and negative radially symmetric solutions of semilinear elliptic equations of the form Δu + F (u, |x|; λ) = 0 in an annulus Ω = {x ∈ R^{n} : A < |x| < B}. The general results are presented and subsequently certain examples are considered.